the ratio of votes for samuel to votes for nicholas in an election is 7:18. 945 people voted for samuel. How many people voted for nicholas?
Answer by Maths68(1474)

Let x people voted for Nicholas
7:18=945:x
7/18=945/x
7x=945*18
7x=17010
7x/7=17010/7
x=2430
2430 people voted for Nicholas